give me a example of vector addition and subtraction?

Grade:11

1 Answers

Kasturi Laturkar
13 years ago

Dear Rushabh,

In both Vector addition and subtraction i,j and k vectors are added seperately .

 

Example for Vector Addition:

(3i+4j+8k) + (5i+2j+7k) = (3+5)i + (4+2)j + (8+7)k= 8i+6j+15k

Example of Vector Subtraction:

(7i+5j+2k) - (2i+3j+5k) = (7-2)i + (5-3)j+ (2-7)k = 5i+2j+(-5)k = 5i+2j-5k